OpenStudy (anonymous):

HELP ME PLZZ WILL MEDAL AND FAN!!!! BUT PLZZ HELP! Evaluate the function x = {-4, -3, -2.5, 3}. x = -4, y = x = -3, y = x = -2.5, y = x = 3, y =

OpenStudy (popitree):

where are the y values?

OpenStudy (popitree):

put -4 in the equation solve for y.. then put -3 solve for y..

OpenStudy (anonymous):

i dont know how to do this can u give me the awenser plz?

OpenStudy (anonymous):

canu do the rest of the Q plz

OpenStudy (popitree):

if u put -2, u will get 11--> 4-(-6)+1 = 11 -4, 16 -(-12)+1 = 29

OpenStudy (anonymous):

did u download the .doc?

OpenStudy (popitree):

i checked that doc, but i never used plotting machine ;) always used to do graph on hand

OpenStudy (anonymous):

kk thank u